Janny Corp. is experiencing rapid growth. Dividends are expected to grow at 30 percent per year during the next three years and then slow down to 8 percent per year, indefinitely. The required rate of return on this stock is 13 percent and the company just paid a $2.40 dividend.

a) What are the expected values of DIV1, DIV2, DIV3, and DIV4?

b) What is the expected stock price three years from now?

c) What is the stock price today? d) Find the dividend yield, DIV1/P0. 3

e) What will next year’s stock price, P1, be?

f) What is the expected rate of return to an investor who buys the stock now and sells it in one year?
Answer #1

D1 = 2.40(1.30) = $3.12

D2 = 2.40(1.30)2 = $4.06

D3 = 2.40(1.30)3 = $5.27

D4 = 5.27(1.08) = $5.69

b.

Stock Price in 3 years = 5.69/(0.13 - 0.08) = $113.80

c.

Stock Price today = 3.12/(1.13) + 4.06/(1.13)2 + (113.80 + 5.27)/(1.13)3

Stock Price today = $88.46

d.

Dividend Yield = 3.12/88.46 = 3.53%